OpenStudy (anonymous):

Explain please: the ratio of Mark's allowance to Jim's allowance is 3:7 The ratio of Mark's allowance to Wanda's allowance is 2:5. What is the ratio of Jim's allowance to Wanda's allowance?

OpenStudy (phi):

you know M/J (it is 3/7) you know M/W (2/5) what is J/W ? one way to do this is "solve" for J and solve for W, then take the ratio

OpenStudy (phi):

if M/J= 3/7 then (flip both) J/M= 7/3 multiply both sides by M to get J= 7M/3 can you solve for W in M/W = 2/5 ?

OpenStudy (anonymous):

Yes, w = 2m/5

OpenStudy (phi):

it looks upside down to me

OpenStudy (anonymous):

Right, sorry. w = 5m/2

OpenStudy (phi):

now we do J/W \[ \frac{\frac{7m}{3}}{\frac{5m}{2}} \]

OpenStudy (anonymous):

so 7m/3:5m/2 7/3:5/2 7/3*2/5 14/15?

OpenStudy (phi):

instead of dividing by 5m/2, multiply by 2/5m

OpenStudy (phi):

yes, that looks good

OpenStudy (anonymous):

Okay thank you.